OpenStudy (anonymous):

Use even odd properties tan(-11π/6)

zepdrix (zepdrix):



zepdrix (zepdrix):

Hmm I'm not sure what they mean by even odd properties. I guess you could consider the fact that tangent is an `odd` function. Meaning that, \[\large \tan\left(-\frac{11\pi}{6}\right) \qquad = \qquad -\tan\left(\frac{11\pi}{6}\right)\]

zepdrix (zepdrix):

If you don't remember your tangent special angles, you can always rewrite it in terms of sines and cosines,\[\large -\tan\left(\frac{11\pi}{6}\right) \qquad = \qquad -\frac{\sin\left(\dfrac{11\pi}{6}\right)}{\cos\left(\dfrac{11\pi}{6}\right)}\]

OpenStudy (anonymous):

Ok so what would the answer be?